40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 20 20:20:36 2020, gautam, Update, PSL, PMC servo checkout PMCsweep.pdfOLTFmeas.pdf
    Reply  Wed Jan 22 11:22:39 2020, gautam, Update, PSL, PMC modulation depth measurement modDepth.pdf
       Reply  Wed Jan 22 20:12:36 2020, gautam, Update, PSL, PMC demodulator electrical characterization demodChar.pdfmixerChar.pdf
          Reply  Thu Jan 23 14:37:05 2020, gautam, Update, PSL, PMC VGA chip damaged? VGAchar.pdf
             Reply  Thu Jan 23 16:37:14 2020, rana, Update, PSL, PMC VGA chip damaged? 
             Reply  Thu Jan 23 22:10:01 2020, gautam, Update, PSL, PMC servo pulled out 
                Reply  Fri Jan 24 15:42:08 2020, gautam, Update, PSL, PMC servo restored 
                   Reply  Sun Jan 26 13:47:00 2020, gautam, Update, PSL, PMC servo characterization elecTFs.pdfVGAchar_postFix.pdfVGAlinearity_postFix.pdfnewOLTFs.pdf
                      Reply  Tue Jan 28 14:33:24 2020, gautam, Update, PSL, Inferred free-running frequency noise inLoopNoise_IMClocked.pdffreqNoiseComparison.pdf
          Reply  Thu Jan 23 18:52:31 2020, gautam, Update, PSL, PMC RFPD characterization PDresp.pdf
Message ID: 15163     Entry time: Tue Jan 28 14:33:24 2020     In reply to: 15156
Author: gautam 
Type: Update 
Category: PSL 
Subject: Inferred free-running frequency noise 

To conclude my PMC noise investigations: Attachment #1 shows the PMC noise inferred from the calibrations earlier in this thread and the fitted OLTF for the PMC loop. Attachment #2 compares the frequency noise (inferred from the error point of the PMC servo) when the IMC is locked / unlocked. I don't know what to make of the fact that the PMC suggests improvement from ~20 Hz onwards already - does this mean that the NPRO noise model is wrong by 1 order of magnitude at 30 Hz?

  • The IMC was locked for the measurement shown in Attachment #1.
  • The in-loop spectra of the error (at the I/F output of the mixer) and control (at TP3) signals were measured with the SR785.
  • The control signal voltage monitors don't seem to work - neither the front panel LEMO nor the signals hooked up to the CDS system show me sensible shapes for the spectra between 1-3 Hz.
  • To convert in loop to free-running, I multiplied the measured error (control) signal spectra by \left | 1-L \right | (\left | \frac{L}{1-L} \right |), where L is the OLTF. THe control signal was pre-processed by multiplying by a pole at 11.3 Hz, corresponding to the LPF formed by the 63.3 kohm series resistor and the 225 nF PZT capacitance.
  • The "NPRO noise model" curve is 10^4/f Hz/rtHz.

While I initially thought the 1/f^2 rise below ~100 Hz is attributable to the IMC cavity length fluctuations, I found that this profile is present even in the measurement with the PSL shutter closed. I am not embarking on a detailed PMC noise budgeting project for now. Note however that we are not shot noise limited anywhere in this measurement band.

  • The measured TF (dots in Attachment #5) was fit with LISO (solid lines in Attachment #5) to allow inferring the out-of-loop servo noise by monitoring the in-loop noise (that plot to follow).
Attachment 1: inLoopNoise_IMClocked.pdf  126 kB  Uploaded Wed Jan 29 11:08:16 2020  | Hide | Hide all | Show all
inLoopNoise_IMClocked.pdf
Attachment 2: freqNoiseComparison.pdf  127 kB  Uploaded Wed Jan 29 11:08:25 2020  | Show | Hide all | Show all
ELOG V3.1.3-